How would I solve this problem?
A person sold 100 shares of a stock at a profit of
13%. If the selling price for the 100 shares was
$2,260, what was paid for the stock?

Answer by Theo(13342)

he sold them for 2260 and made a profit of 13%.

let x = what he paid for them.

then x + .13 * x = 2260

factor out the x to get:

(1 + .13) * x = 2260

simplify to get:

1.13 * x = 2260

divide both sides of this equation by 1.13 to get:

x = 2260 / 1.13 = 2000